# Towards a General-Purpose Zero-Shot Synthetic Low-Light Image and Video Pipeline

### v1

@@ 4.5 Ablation Study
We highlight the importance of our MLP head, demonstrating a major performance increase when added to the U-Net, with a substantial improvement of 382.916 in FID score between model v1 and model v2, and therefore it is a crucial component for allowing our network to mimic the noise characteristics accurately. Furthermore, we verify that our $\mathcal{L}_{rec}$ improves our network performance by assisting in understanding noise in an image/video. This ablation study confirms the significance of each component in our proposed DEN. The results are shown in Table II.
TABLE II: Ablation Study of the DEN Architecture.
@@ 5 Conclusion
In our future work, we will focus on video blur, as most existing synthetic pipelines focus on image-only degradations. We would also like to address more complex noise types, such as spatially-correlated noise due to compression, and noise from in-camera processing. Future work could also explore the use of unsupervised algorithms for the network to group similar noise characteristics, to synthesize camera-specific noise patterns without explicit camera information. Furthermore, our model currently focuses on noise synthesis in low-light videos. We would like to expand our method to also map the illumination of the video from normal-light to low-light in a realistic manner, allowing easier low-light synthesis for downstream tasks.
Acknowledgments
This work was supported by the UKRI MyWorld Strength in Places Programme (SIPF00006/1) and EPSRC Doctoral Training Partnerships (EP/W524414/1)

### v2

@@ 4.5. Ablation Study
We highlight the importance of our MLP head, demonstrating a major performance increase when added to the U-Net, with a substantial improvement of 382.916 in FID score between model v1 and model v2, and therefore it is a crucial component for allowing our network to mimic the noise characteristics accurately. Furthermore, we verify that our $\mathcal{L}_{rec}$ improves our network performance by assisting in understanding noise in an image/video; the added decoder head and accompanying $\mathcal{L}_{rec}$, forces the model to learn the specifics of the reference noise, as seen in the increase across all metrics for model v3. This ablation study confirms the significance of each component in our proposed DEN.
Table 2. Ablation Study of the DEN Architecture.
@@ 5. Conclusion
In our future work, we will explore video-specific degradations, such as motion blur, as most existing synthetic pipelines focus on image-only degradations. We would like to also address more complex degradations, like spatially-correlated noise, arising from the camera ISP. Furthermore, our model currently focuses on noise synthesis in low-light videos; we would like to expand our method to learn the illumination mapping and color distortions as well, allowing easier low-light synthesis for downstream tasks. Finally, we aim to explore using our method for joint low-light enhancement and synthesis by improving DEN’s reconstruction capabilities.
Acknowledgements.
This work was supported by the UKRI MyWorld Strength in Places Programme (SIPF00006/1), the EPSRC Doctoral Training Partnerships (EP/W524414/1), and an EPSRC iCASE grant in collaboration with the British Broadcasting Corporation (BBC).